Do the calculations to make 200ml of 0.1M sodium acetate buffer with a pH of 5.1 starting with 5.0M acetic acid and 1.0M NaOH ****show calculations***** 1.How much acetic acid is needed? 2.What is the ratio of Ac- to HAc at pH of 5.1? 3. What fraction of total acetate is Ac-at pH 5.1 4. How much OH- is needed to obtain 68.7% Ac-?
Answer #1

From tables:

pK_a=4.75

Using the Henderson-Hasselbach equation:

pH=pK_a+Log\frac{[Ac^-]}{[HAc]}

Then:

5.1=4.75+Log\frac{[Ac^-]}{[HAc]}

\frac{[Ac^-]}{[HAc]}=2.24

\frac{n_{Ac^-}}{n_{HAc}}=2.24

\frac{V_{NaOH}*1}{V_{HAc}*5-V_{NaOH}*1}=2.24

3.24V_{NaOH}=11.2V_{HAc}

3.24(0.2-V_{HAc})=11.2V_{HAc}

(0.2-V_{HAc})=3.46V_{HAc}

V_{HAc}=0.045L=45mL

V_{NaOH}=155mL